# NuwaTS: a Foundation Model Mending Every Incomplete Time Series

### v2

@@ D.1 Learning From Incomplete Training data
D.2 The effect of the number of the GPT2 layer.
@@
D.3 Effect of the Type of PLM Backbone
Table 11: Effect of different PLM backbone. A lower MSE or MAE indicates a better imputation performance. Red: the best.
@@ D.3 Effect of the Type of PLM Backbone
We experimented with three different backbones in Table 11: GPT2 [30], BERT [7], and LLaMA [35], uniformly using the first six layers of each. Among them, BERT is a transformer with bi-directional attention, while the other two, GPT2 and LLaMA, employ causal attention.
@@
Based on the BERT model pre-trained on general fused dataset, We have $\mathbf{E}_{i}=[\mathbf{k},\mathbf{z}_{i,(v_{g})},\mathbf{E}_{i,(p)}]$ before feeding into the backbone, where $\mathbf{E}_{i}\in\mathbb{R}^{D\times(N+2)}$. We feed the $\mathbf{E}_{i}$ into the backbone, and get the final representation $\mathbf{h}^{(\text{Layer})}_{i}\in\mathbb{R}^{(N+2)\times D}$. Then We extract the first embedding $\mathbf{k^{\prime}}\in\mathbb{R}^{D}$ from $\mathbf{h}^{(\text{Layer})}_{i}$ and take it as [CLS] token [7] which is mixed with the patch embeddings through bidirectional attention mechanism. With same approach as Appendix C.3, we obtain scattered points $\mathbf{k^{\prime}}^{p}\in\mathbb{R}^{N\times 2}$ and visualize them in Figure 13.
Figure 13: Domain information visualization. We take the model’s first output embedding as the [CLS] token, which carries domain information from following sequences. We then use the t-SNE [37] method to visualize the domain information from different datasets.
D.4 Embed with simple linear layer VS. Embed with text-alignment.
For tokenization, we employ a simple linear layer to embed patches. We acknowledge that recent research [12] has proposed utilizing Patch Reprogramming mechanism to align time patches with PLM’s pre-trained word embeddings, thereby activating the model’s time series understanding and reasoning capabilities. However, we contend that, due to the high proportion of missing values in the input masked patches and the varying locations of these missing values, modality alignment is not effective in representing such complex incomplete time series. Table 12 shows that simple linear embedding is better than the text-alignment strategy.
Table 12: Embedding methods comparision.
@@
Embed with text-alignment [12] 0.250 0.357 0.024 0.100 0.128 0.242 0.014 0.076 0.205 0.311 0.323 0.153
@@ Appendix E Main Results in Detailed
Table 13: Detailed Imputation results on ETTs, ECL and weather with missing rate set to 0.1. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 14: Detailed Imputation results on ETTs, ECL and weather with missing rate set to 0.2. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 15: Detailed Imputation results on ETTs, ECL and weather with missing rate set to 0.3. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 16: Detailed Imputation results on ETTs, ECL and weather with missing rate set to 0.4. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
[b] Model ETTh1 ETTh2 ETTm1 ETTm2 ECL Weather PEMS03 PEMS04 PEMS07 PEMS08 MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Median 0.710 0.606 0.716 0.469 0.687 0.579 0.736 0.461 0.988 0.827 0.984 0.493 0.688 0.607 0.736 0.642 0.749 0.643 0.728 0.650 Last 0.411 0.466 0.073 0.139 0.318 0.403 0.045 0.113 0.930 0.816 0.701 0.343 0.455 0.494 0.468 0.506 0.488 0.506 0.428 0.484 Autoformer(2021) 0.426 0.487 0.347 0.367 0.238 0.349 0.113 0.239 0.103 0.232 0.501 0.393 0.419 0.500 0.231 0.374 0.213 0.356 0.632 0.608 Fedformer(2022) 0.280 0.395 0.447 0.421 0.053 0.164 0.033 0.129 0.100 0.231 0.174 0.163 0.171 0.310 0.255 0.378 0.217 0.350 0.199 0.321 Dlinear(2023) 0.194 0.311 0.046 0.138 0.084 0.200 0.055 0.167 0.144 0.275 0.182 0.149 0.073 0.200 0.080 0.207 0.064 0.185 0.073 0.195 iTransformer(2024) 0.565 0.542 0.230 0.309 0.235 0.345 0.213 0.341 0.069 0.175 0.412 0.357 0.094 0.221 0.110 0.240 0.083 0.208 0.135 0.266 BRITS(2018) 0.162 0.273 0.086 0.147 0.062 0.145 0.047 0.101 0.270 0.395 0.781 0.472 0.140 0.264 0.253 0.364 0.224 0.350 0.217 0.336 TimesNet(2022) 0.128 0.249 0.017 0.085 0.041 0.128 0.009 0.059 0.336 0.433 0.664 0.274 0.093 0.210 0.144 0.259 0.123 0.242 0.106 0.221 PatchTST(2023) 0.148 0.269 0.019 0.089 0.061 0.163 0.009 0.062 0.093 0.227 0.218 0.11 0.050 0.160 0.064 0.180 0.044 0.148 0.058 0.169 SAITS(2023) 0.150 0.257 0.191 0.174 0.066 0.139 0.132 0.126 0.429 0.503 0.936 0.486 0.155 0.284 0.267 0.377 0.226 0.347 0.246 0.361 GPT4TS(2024) 0.143 0.252 0.021 0.085 0.045 0.128 0.010 0.058 0.258 0.375 0.945 0.317 0.090 0.208 0.149 0.262 0.122 0.240 0.099 0.214 NuwaTS(specific) 0.145 0.265 0.017 0.086 0.050 0.140 0.009 0.059 0.058 0.160 0.264 0.117 0.041 0.139 0.050 0.149 0.033 0.119 0.049 0.145 PatchTST(one-for-all) 0.136 0.254 0.016 0.082 0.051 0.147 0.009 0.060 0.085 0.212 0.187 0.104 0.048 0.159 0.058 0.171 0.043 0.148 0.052 0.159 NuwaTS(one-for-all) 0.119 0.225 0.015 0.077 0.040 0.118 0.009 0.055 0.057 0.157 0.162 0.075 0.039 0.135 0.050 0.148 0.033 0.119 0.044 0.136 NuwaTS(fine-tuned) 0.115 0.220 0.014 0.076 0.040 0.117 0.008 0.054 0.055 0.155 0.168 0.074 0.039 0.135 0.050 0.148 0.033 0.119 0.044 0.136
Table 17: Detailed Imputation results on ETTs, ECL and weather with missing rate set to 0.5. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 18: Detailed Imputation results on ETTs, ECL and weather with missing rate set to 0.6. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 19: Detailed Imputation results on ETTs, ECL and weather with missing rate set to 0.7. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 20: Detailed Imputation results on ETTs, ECL and weather with missing rate set to 0.8. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 21: Detailed Imputation results on ETTs, ECL and weather with missing rate set to 0.9. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@ Appendix F Limitation
The model was trained on time series segments fixed at a length of 96, imputing the masked series based on the unmasked portions. Although our model can adapt to different domains and missing patterns, showing strong adaptability after domain-transfer fine-tuning, in practical applications, the model may require further fine-tuning when dealing with segments longer than 96 as well as the segments that entirely missed. Additionally, NuwaTS discards multivariate correlations, relying on temporal information and identified domain information, which may degrade the imputation performance. For instance, if a sensor loses a large amount of data, the data from adjacent sensors can be used as supplementary information. Future work will aim to improve NuwaTS’s capability in handling longer missing segments and incorporate multivariate correlations.
Appendix G Broader Impact
We believe that our NuwaTS model not only plays an important role in the time series domain but also that its one-for-all training paradigm and domain-specific fine-tuning method will inspire broader applications in the field of data mining. In practical applications, NuwaTS can be widely used to address various time series data gaps, and the imputed data can enhance the performance of downstream tasks. However, there may be some discrepancies between the time series imputed by NuwaTS and the real data. Therefore, it is necessary to validate the improvements in downstream tasks through multiple methods. Additionally, NuwaTS is not recommended for use in investment-related forecasting, such as stock trading.

### v3

@@ D.3 Learning From Incomplete Training data
D.4 The effect of the number of the GPT2 layer.
@@
D.5 Effect of the Type of PLM Backbone
Table 13: Effect of different PLM backbone. A lower MSE or MAE indicates a better imputation performance.  Green : the best.
@@ D.5 Effect of the Type of PLM Backbone
We experimented with three different backbones in Table 13: GPT2 (Radford et al., 2019), BERT (Devlin et al., 2019), and LLaMA (Touvron et al., 2023), uniformly using the first six layers of each. Among them, BERT is a transformer with bi-directional attention, while the other two, GPT2 and LLaMA, employ causal attention.
@@
Based on the BERT model pre-trained on general fused dataset, We have $\mathbf{E}_{i}=[\mathbf{k},\mathbf{z}_{i,(v_{g})},\mathbf{E}_{i,(p)}]$ before feeding into the backbone, where $\mathbf{E}_{i}\in\mathbb{R}^{D\times(N+2)}$. We feed the $\mathbf{E}_{i}$ into the backbone, and get the final representation $\mathbf{h}^{(\text{Layer})}_{i}\in\mathbb{R}^{(N+2)\times D}$. Then We extract the first embedding $\mathbf{k^{\prime}}\in\mathbb{R}^{D}$ from $\mathbf{h}^{(\text{Layer})}_{i}$ and take it as [CLS] token (Devlin et al., 2019) which is mixed with the patch embeddings through bidirectional attention mechanism. With same approach as Appendix C.3, we obtain scattered points $\mathbf{k^{\prime}}^{p}\in\mathbb{R}^{N\times 2}$ and visualize them in Figure 22.
Figure 22: Domain information visualization. We take the model’s first output embedding as the [CLS] token, which carries domain information from following sequences. We then use the t-SNE (Van der Maaten & Hinton, 2008) method to visualize the domain information from different datasets.
D.6 Embed with simple linear layer VS. Embed with text-alignment.
For tokenization, we employ a simple linear layer to embed patches. We acknowledge that recent research (Jin et al., 2024) has proposed utilizing Patch Reprogramming mechanism to align time patches with PLM’s pre-trained word embeddings, thereby activating the model’s time series understanding and reasoning capabilities. However, we contend that, due to the high proportion of missing values in the input masked patches and the varying locations of these missing values, modality alignment is not effective in representing such complex incomplete time series. Table 14 shows that simple linear embedding is better than the text-alignment strategy.
Table 14: Embedding methods comparison.
@@
Embed with text-alignment (Jin et al., 2024) 0.250 0.357 0.024 0.100 0.128 0.242 0.014 0.076 0.205 0.311 0.323 0.153
@@ Appendix E Main Results in Detailed
Table 15: Detailed Imputation results with missing rate set to 0.1. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 16: Detailed Imputation results with missing rate set to 0.2. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 17: Detailed Imputation results with missing rate set to 0.3. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 18: Detailed Imputation results with missing rate set to 0.4. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
[b] Model ETTh1 ETTh2 ETTm1 ETTm2 ECL Weather PEMS03 PEMS04 PEMS07 PEMS08 MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Median 0.710 0.606 0.716 0.469 0.687 0.579 0.736 0.461 0.988 0.827 0.984 0.493 0.688 0.607 0.736 0.642 0.749 0.643 0.728 0.650 Last 0.411 0.466 0.073 0.139 0.318 0.403 0.045 0.113 0.930 0.816 0.701 0.343 0.455 0.494 0.468 0.506 0.488 0.506 0.428 0.484 Autoformer(2021) 0.426 0.487 0.347 0.367 0.238 0.349 0.113 0.239 0.103 0.232 0.501 0.393 0.419 0.500 0.231 0.374 0.213 0.356 0.632 0.608 Fedformer(2022) 0.280 0.395 0.447 0.421 0.053 0.164 0.033 0.129 0.100 0.231 0.174 0.163 0.171 0.310 0.255 0.378 0.217 0.350 0.199 0.321 Dlinear(2023) 0.194 0.311 0.046 0.138 0.084 0.200 0.055 0.167 0.144 0.275 0.182 0.149 0.073 0.200 0.080 0.207 0.064 0.185 0.073 0.195 iTransformer(2024) 0.565 0.542 0.230 0.309 0.235 0.345 0.213 0.341 0.069 0.175 0.412 0.357 0.094 0.221 0.110 0.240 0.083 0.208 0.135 0.266 BRITS(2018) 0.162 0.273 0.086 0.147 0.062 0.145 0.047 0.101 0.270 0.395 0.781 0.472 0.140 0.264 0.253 0.364 0.224 0.350 0.217 0.336 TimesNet(2022) 0.128 0.249 0.017 0.085 0.041 0.128 0.009 0.059 0.336 0.433 0.664 0.274 0.093 0.210 0.144 0.259 0.123 0.242 0.106 0.221 PatchTST(2023) 0.148 0.269 0.019 0.089 0.061 0.163 0.009 0.062 0.093 0.227 0.218 0.110 0.050 0.160 0.064 0.180 0.044 0.148 0.058 0.169 SAITS(2023) 0.150 0.257 0.191 0.174 0.066 0.139 0.132 0.126 0.429 0.503 0.936 0.486 0.155 0.284 0.267 0.377 0.226 0.347 0.246 0.361 GPT4TS(2024) 0.143 0.252 0.021 0.085 0.045 0.128 0.010 0.058 0.258 0.375 0.945 0.317 0.090 0.208 0.149 0.262 0.122 0.240 0.099 0.214 NuwaTS(specific) 0.145 0.265 0.017 0.086 0.050 0.140 0.009 0.059 0.058 0.160 0.264 0.117 0.041 0.139 0.050 0.149 0.033 0.119 0.049 0.145 PatchTST(one-for-all) 0.136 0.254 0.016 0.082 0.051 0.147 0.009 0.060 0.085 0.212 0.187 0.104 0.048 0.159 0.058 0.171 0.043 0.148 0.052 0.159 NuwaTS(one-for-all) 0.119 0.225 0.015 0.077 0.040 0.118 0.009 0.055 0.057 0.157 0.162 0.075 0.039 0.135 0.050 0.148 0.033 0.119 0.044 0.136 NuwaTS(fine-tuned) 0.115 0.220 0.014 0.076 0.040 0.117 0.008 0.054 0.055 0.155 0.168 0.074 0.039 0.135 0.050 0.148 0.033 0.119 0.044 0.136
Table 19: Detailed Imputation results with missing rate set to 0.5. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 20: Detailed Imputation results with missing rate set to 0.6. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 21: Detailed Imputation results with missing rate set to 0.7. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 22: Detailed Imputation results with missing rate set to 0.8. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@
Table 23: Detailed Imputation results with missing rate set to 0.9. We set the input length to 96. A lower MSE or MAE indicates a better imputation performance.
@@ Appendix F Limitation
The model was trained on time series segments fixed at a length of 96, imputing the masked series based on the unmasked portions. Although our model can adapt to different domains and missing patterns, showing strong adaptability after domain-transfer fine-tuning, in practical applications, the model may require further fine-tuning when dealing with segments longer than 96, as well as the segments that are entirely missed.
Future work will aim to improve NuwaTS’s capability in handling longer missing segments.